# OmniSVG: A Unified Scalable Vector Graphics Generation Model

### v1

@@ 1 Introduction
To generate SVG contents, existing methods adopt either optimization-based methods or auto-regressive approaches. Optimization-based methods [19, 30, 56] iteratively refine the SVG parameters by optimizing differentiable vector graphics rasterizers. Though these methods are effective in generating SVG icons, they suffer from significant computational overhead when scaling up to intricate samples and produce unstructured outputs with redundant anchor points. In contrast, auto-regressive methods build transformer models or adapt pre-trained Large Language Models (LLMs) to directly generate XML parameters [55] or codes [53, 37] to represent SVGs. Benefiting from an end-to-end learning pipeline, the auto-regressive method is a more scalable approach [5] as it is able to learn directly from a large collection of SVG samples. However, existing auto-regressive approaches are limited to basic SVG contents because of their limited context window length and the scarcity of complex SVG data [9, 20, 49]. To address these limitations, we propose OmniSVG, the first unified framework that harnesses pre-trained VLMs for end-to-end multimodal complex SVG generations. By parameterizing SVG coordinates and commands into discrete tokens, OmniSVG  decouples structural logic from low-level geometry, mitigating the “coordinate hallucination” problem prevalent in code-based LLMs, and produces vivid and colorful SVG results. Benefiting from the next token prediction training objective, OmniSVG  is also capable of completing SVGs with diverse generation results given some partial observations. Compared to traditional auto-regressive SVG generation, our method is able to parameterize the SVGs with a token length up to $30k$, facilitating the generation of complex and high-quality SVGs. Building upon pre-trained VLMs, our method natively integrates the ability to reason upon the visual and textual instructions to synthesize editable, high-fidelity SVGs across diverse domains, from icons to intricate illustrations and anime characters.
To advance the development of SVG synthesis, we introduce MMSVG-2M, a multi-modal SVG synthesis dataset with two million richly annotated assets, encompassing icons, illustrations, and anime designs. We also establish a standardized evaluation protocol, MMSVG-Bench, for three key SVG generation tasks, i.e. “Text-to-SVG”, “Image-to-SVG”, and “Character Reference SVG Generation”. Extensive experiments show that OmniSVG  produces highly detailed and complex SVG contents, surpassing prior arts both quantitatively and qualitatively.
@@
  • We introduce OmniSVG, the first family of end-to-end multimodal SVG generators that leverage pre-trained Vision-Language Models (VLMs), capable of generating complex and detailed SVGs, from simple icons to intricate anime characters.
  • We present MMSVG-2M, a large-scale dataset comprising two million SVG assets, along with a standardized evaluation protocol for various conditional SVG generation tasks, providing a comprehensive resource for future research.
@@ 2 Related Works
Image Vectorization. Recent advances in vectorization leverage diffusion models paired with differentiable rasterizers, such as DiffVG [25]. These methods aim to convert raster images into SVG paths using score distillation sampling (SDS) [32]. Specifically, LIVE [30] employs a layer-wise optimization strategy with self-crossing loss and undirected distance-guided focus loss; VectorFusion [19] integrates a pre-trained text-to-image diffusion model (Stable Diffusion [38]) with SVG-specific regularizers; and SVGDreamer [7] introduces a semantic-driven image vectorization process and particle-based SDS loss to enhance visual quality. While these methods achieve remarkable results, they face limitations such as over-smoothing, color over-saturation, and lack of editability, often producing tangled paths that fail to capture hierarchical structures inherent in professional SVG designs.
SVG Generation. Early attempts to generating SVGs directly builds sequence models like RNNs [15, 36, 16, 40, 41], VAEs [4, 28, 45, 42, 47], and Transformers [4, 52] to compress SVG commands into latent representations. Meanwhile, DeepSVG [4] further parameterizes SVGs using a dual transformer architecture but struggles with geometric consistency. Recently, the advent of large language models (LLMs) [26, 60, 48, 57, 5, 6, 59, 58] unleashes the potential of generating SVGs via XML code synthesis. StrokeNUWA [44] converts vector images into a compressed representation as a sequence of tokens with VQVAEs. LLM4SVG [55] employs a structured SVG encoding approach, utilizing learnable semantic tokens to accurately represent SVG components and their properties. StarVector [37] combines an image encoder with an LLM to generate SVG code from images, but this approach separates the understanding of text and images, leading to a lack of effective alignment between textual descriptions and visual outputs. However, these methods suffer from a limited token context window, which may struggle to process highly complex SVGs that require longer sequences. However, these methods [53, 37, 55] still face significant limitations when handling complex SVGs, as illustrations and anime characters often reach a token length up to $10k$ or even $30k$.
@@
SVG Datasets and Benchmarks. The lack of suitable datasets for complex SVG structures presents a significant challenge. Existing datasets [9, 20, 49] primarily focus on simplified path-based SVGs or monochrome icons, overlooking the intricate layered structures and rich color semantics found in real-world designs. For instance, the FIGR-8-SVG [9] dataset is centered on monochromatic icons, while StarVector [37] proposes categorized datasets, including illustrations, icons, emojis, and fonts. However, its datasets only include examples with up to $8.2k$ tokens, which still fail to capture the complexities of layered structures and rich color semantics. Benchmark evaluations, such as VGBench [65], further highlight gaps in multi-format testing and the absence of comprehensive coverage for illustrative SVGs.
To address these shortcomings, we introduce MMSVG-2M, a multimodal SVG synthesis dataset comprising two million richly annotated assets, including icons, illustrations, and complex anime designs. Additionally, we present a standardized evaluation benchmark, MMSVG-Bench, for SVG generation tasks of varying complexity. Unlike prior efforts, our benchmark emphasizes real-world applicability, offering full public access and rigorous evaluation across key metrics such as generation fidelity, diversity, and editability. This foundation lays the groundwork for future research, enabling more versatile and user-centric SVG synthesis.
@@ 3.1 Data Curation
3.2 Multimodal Instruction Following for SVG
MMSVG-Bench mainly focuses on the following SVG generation tasks, including:
Text-to-SVG requires a model to generate SVGs from text instructions. This includes the MMSVG-2M dataset’s test set, which provides textual description per image, and the FIGR-8-SVG dataset, enabling the generation of simpler (fewer paths) icons from text. For text-to-SVG evaluation, we measure visual quality with Frechet Inception Distance (FID) [46], text-SVG alignment with CLIP score [34], aesthetic appeal with Aesthetic score [39], and HPS (Human Preference Scores) [54].
Image-to-SVG evaluates a model’s ability to convert images into SVGs. MMSVG-Icon, MMSVG-Illustration, and MMSVG-Character provide increasing levels of visual complexity. We provide train, validation, and test splits for these datasets. For image-to-SVG evaluations, we use DinoScore (DINO) [31], the cosine similarity between DinoV2 features, Structural Similarity Index (SSIM) [50], Learned Perceptual Image Patch Similarity (LPIPS) [61], and Mean Squared Error (MSE) to evaluate the image-conditioned SVG generation.
Character-Reference SVG Generation evaluates a model’s ability to generate novel SVGs while keeping the profile of the characters depicted in the input image. The difference between the image-to-SVG task and the character-reference SVG Generation task is that the character-reference SVG generation task is not to reconstruct the SVG of the input image, but to recreate a specific character SVG for the input image, as shown in Fig. 6. Acknowledging the absence of established baselines, we assess the alignment between input character images and generated SVGs using the third-party VLMs assistant GPT-4o [18] . This assistant assigns a score ranging from 1 to 10, where 1 indicates a poor match with significant discrepancies, and 10 signifies an outstanding alignment with a highly accurate representation of the original image.
For evaluation metrics, we also calculate the average token length (# tokens) of a generated SVG sample utilizing the Qwen2.5-VL [1] tokenizer. For computational cost, we compute the average generation time of generating per SVG sample.
@@ 4 OmniSVG
SVG Parameterization. As illustrated in Sec. 3, our MMSVG-2M dataset simplifies every SVG by removing all attributes and using five basic commands, including “Move To” (M), “Line To” (L), “Cubic Bézier” (C), “Elliptical Arc” (A), “ClosePath” (Z). After the simplification, an SVG script $G$ is represented as the combination of $M$ paths, $G=\{P_{i}\}_{i=1}^{M}$. Here, $P_{i}$ is the $i$-th path containing $N_{i}$ commands, $P_{i}=\{C_{i}^{j}\}_{j=1}^{N_{i}}$, where $C_{i}^{j}$ is the $j$-th command in the $i$-th path. Each command is represented as $C_{i}^{j}=(U_{i}^{j},V_{i}^{j})$, containing both the type identifier $U_{i}^{j}\in\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z}\}$ and the corresponding location argument $V_{i}^{j}$. “Move To” (M) command means to move the cursor to the end-point $(x_{2},y_{2})$ without drawing anything. “Line To” (L) means to draw a line to the point $(x_{2},y_{2})$. “Cubic Bézier” (C) means to draw a cubic Bézier curve with control points $(q_{x1},q_{y1})$, $(q_{x2},q_{y2})$ and end-point $(x_{2},y_{2})$. “Elliptical Arc” (A) means to draw an elliptical arc with radii $r_{x}$ and $r_{y}$ (semi-major and semi-minor axes), rotated by angle $\varphi$ to the x-axis, and end-point ($x_{2}$, $y_{2}$). And the “ClosePath” (Z) means to close the path. The visualization of the commands is shown in Tab. 2. However, these commands are only used to draw line strokes without color. To fill the gap in the SVG fill attribute gap, we assign distinctive numbers to the hex form of the SVG Fill ($\mathrm{F}$) attribute, distinguishing it from the original SVG commands and coordinates. This command means to fill the color of each path, which contains essential SVG commands and coordinates. To this end, we are able to use a total six types of commands $U_{i}^{j}\in\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z},\mathrm{F}\}$ to parameterize a colored SVG parameterization.
Model Architecture. OmniSVG leverages the Qwen2.5-VL architecture, a vision-language models (VLMs), which excels in processing both visual and textual inputs, enabling the generation of precise and compact SVG outputs. OmniSVG is trained for text-to-SVG, image-to-SVG, and character-reference SVG generation task.
As depicted in Fig. 2, We first tokenize and embed the interleaved text and image inputs as prefix tokens. Then, we tokenize the SVG scripts into a sequence and concatenate it to the end of the prefix token with the SVG tokenizer. The full sequence is then served as the input of the decoder-only language model.
@@
Specifically, our SVG tokenizer transforms SVG scripts $X_{s}$ into an ordered SVG token sequence within the same representation space as the pre-trained VLM. Following IconShop [52], we flatten the layered structure of the SVG script by concatenating different paths into a single command sequence, where each path begins with the drawing commands followed by point coordinates. Therefore, each SVG sequence could be represented as a flattened sequence. As the generation identifier, we apply special tokens like <SOP> and <EOS> to the two ends of a SVG sequence, identifying the begining and ending of a SVG sequence. We assign special tokens for each command type, i.e. $\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z},\mathrm{F}\}$. To shorten the length of the SVG sequence, we further merge the point coordinates into one token with a mapping function: $<x,y>\rightarrow x\times w+y$, where $w$ is the width of the image. The SVG sequence are then lifted into the same embedding space as the pre-trained VLM with a learnable embedding layer.
@@ 6 Conclusions
Limitations and Future Work. During inference, OmniSVG generates tens of thousands of tokens for complex samples, which inevitably leads to a considerable generation time. OmniSVG is only bounded by vector style image prompt and fails on natural images as illustrated in Fig. LABEL:fig:boundary. As for future work, recent endeavors on multi-token prediction [13, 2] and KV-cache compression [63, 3] provide a promising way to save the generation cost. Additionally, the auto-regressive nature of OmniSVG also unlocks future opportunities for in-context learning [62, 64, 43], chain-of-thought reasoning [51, 14], and multi-turn interleaved generation [17, 27], thereby providing a more dexterous user control.
@@ A.2 SVG-Image-Text Pairs Construction
Our MMSVG-2M dataset comprises two million SVG samples with the corresponding rasterized images. We generate captions on the rasterized images with BLIP-2 [24], thereby providing textual descriptions that enable us to fine-tune our model to follow these instructions. We use CairoSVG [21] for rasterization and remove samples that produced completely white images.
@@ A.3 Character-SVG Pairs Construction
As illustrated in the Fig. 1, part of our proposed MMSVG-2M-Character subset is constructed using a generative pipeline. As shown in the pipeline diagram in Fig. 2, we employ a FLUX [22]-based generative model enhanced with a vector-style LoRA to enable the generation of SVG-style data. For image-based conditioning, we adopt FLUX-Redux [23], which injects image features via a SigLIP encoder and projects them into image embeddings. These embeddings are then concatenated with the text tokens as conditioning inputs for FLUX [22]. However, in practice, the original Redux [23] conditioning proves to be overly strong. To address this, we adopt a community-implemented variant of Redux that downsamples the image embeddings in 2D space. As observed in our experiments shown in Fig. 11, a downsampling factor between 2×2$\times$2 × and 3×3$\times$3 × yields the most reasonable SVG-style character references. Finally, we employ VTracer [10] to perform near-instant vectorization of the generated images. To construct the MMSVG-2M-Character subset, we first filter $103k$ character instances from the Danbooru [11] dataset and apply the aforementioned pipeline. We compare the raw FLUX [22] outputs and their vectorized counterparts, retaining only those samples with PSNR and SSIM scores above a certain threshold as valid data.

### v2

@@ 1 Introduction
Existing methods adopt either optimization-based methods or auto-regressive approaches to generate SVG contents. The optimization-based methods [21, 33, 59] iteratively refine the SVG parameters by optimizing differentiable vector graphics rasterizers. Though these methods are effective in generating SVG icons, they suffer from significant computational overhead when scaling up to intricate samples and produce unstructured outputs with redundant anchor points. In contrast, auto-regressive methods build transformer models or adapt pre-trained Large Language Models (LLMs) to directly generate XML parameters [58] or codes [56, 41] representing SVGs. Benefiting from the end-to-end learning pipeline, the auto-regressive method is a more scalable approach [5] as it can learn directly from a large collection of SVG samples. However, existing auto-regressive approaches are limited to basic SVG contents [11, 22, 52] because of the limited context length and the scarcity of complex SVG data.
In this paper, we propose OmniSVG that harnesses native VLMs [1] for various end-to-end multimodal SVG generation tasks. By parameterizing SVG coordinates and commands into discrete tokens, OmniSVG decouples structural logic from low-level geometry, mitigating the “coordinate hallucination” problem prevalent in code-based LLMs, and produces vivid and colorful SVG results. Additionally, the next token prediction training objective enables OmniSVG to complete SVGs with diverse generation results given some partial observations. Compared to traditional auto-regressive SVG generation methods, OmniSVG is able to parameterize SVGs exceeding $30k$ tokens, facilitating the generation of detailed and complex SVG contents. Building upon pre-trained VLMs, our method natively integrates the ability to reason upon visual and textual instructions to synthesize editable, high-fidelity SVGs across diverse domains, from icons to intricate illustrations and anime characters.
To advance the development of SVG synthesis, we introduce MMSVG-2M, a multi-modal SVG synthesis dataset with two million richly annotated assets, encompassing icons, illustrations, and anime designs. We also establish a standardized evaluation protocol, MMSVG-Bench, for three key SVG generation tasks, i.e. “Text-to-SVG”, “Image-to-SVG”, and “Character Reference SVG Generation”. Extensive experiments show that OmniSVG produces highly detailed and complex SVG contents, surpassing prior art both quantitatively and qualitatively.
@@
  • We introduce OmniSVG, a family of end-to-end multimodal SVG generators that leverage native VLMs for generating complex and detailed SVGs, from simple icons to intricate anime characters.
  • We present MMSVG-2M, a large-scale dataset comprising two million SVG assets, along with a standardized evaluation protocol for various multi-modal SVG generation tasks, providing a comprehensive resource for future research.
@@ 2 Related Works
SVG Generation. Early attempts to generating SVGs directly utilize architectures like RNNs [17, 40, 18, 43, 44], VAEs [4, 31, 48, 45, 50, 65], and Transformers [4, 55] to compress SVG commands into latent representations. Meanwhile, DeepSVG [4] further parameterizes SVGs using a dual transformer architecture but struggles with geometric consistency. Recently, the advent of large language models (LLMs) [29, 63, 51, 60, 5, 6, 62, 61, 47] unleashes the potential of generating SVGs via XML code synthesis [58, 56, 41]. However, the limited context length of existing LLM-based SVG generation methods [56, 41, 58] poses significant challenges in handling complex SVGs that exceed $10k$ tokens. In this paper, we explore the potential of native Vision-Language Models (VLMs) in multi-modal SVG generation. Combining pre-trained VLMs with SVG command parameterization, we validate that OmniSVG is able to follow multi-modal instructions and generate more complex SVGs than before.
@@
Image Vectorization. Recent advancements in vectorization harness diffusion models paired with differentiable rasterizers, using techniques like score distillation sampling [36, 21, 7] and specialized regularizers [27, 33] to convert raster images into SVG paths. While these methods achieve remarkable results, they face limitations such as over-smoothing, color over-saturation, and lack of editability, often producing tangled paths that fail to capture hierarchical structures inherent in professional SVG designs. In this paper, we present an end-to-end generation method that directly follows multi-modal instructions to generate high-quality SVGs with improved path clarity.
@@
SVG Datasets and Benchmarks. The lack of suitable datasets for complex SVG structures presents a significant challenge. Existing datasets [11, 22, 52] primarily focus on simplified path-based SVGs or monochrome icons, overlooking the intricate layered structures and rich color semantics found in real-world designs. For example, FIGR-8-SVG [11] focuses on monochromatic icons, while StarVector [41] proposes categorized datasets, including illustrations, icons, emojis, and fonts. Therefore, existing datasets only present simple SVG samples that do not exceed $8.2k$ tokens, failing to capture the complexities of layered structures and rich color semantics. Benchmark evaluations, such as VGBench [71], further highlight gaps in multi-format testing and the absence of comprehensive coverage for illustrative SVGs. To this end, we introduce MMSVG-2M, a multimodal SVG synthesis dataset comprising two million richly annotated assets, including icons, illustrations, and complex anime designs. We also present a standardized evaluation protocol, MMSVG-Bench, to evaluate diverse multi-modal SVG generation tasks with varying complexity.
@@ 3.1 MMSVG-2M
3.2 MMSVG-Bench
As a multi-modal SVG generation evaluation protocol, MMSVG-Bench put emphasis on both the generation quality and how well the model follows multi-modal instructions.
Text-to-SVG requires a model to generate SVGs from text instructions. We measure the visual quality with Frechet Inception Distance (FID) [49], aesthetic appeal with Aesthetic score [42], text-SVG alignment with CLIP score [38], and Human Preference Scores (HPS) [57].
Image-to-SVG evaluates a model’s ability to convert images into SVGs. To quantify the distance between the input and output SVG, we calculate the cosine similarity of DinoV2 features (DinoScore) [34], Structural Similarity Index (SSIM) [53], Learned Perceptual Image Patch Similarity (LPIPS) [66], and Mean Squared Error (MSE).
Character-Reference SVG Generation evaluates a model’s ability to generate novel SVGs while keeping the profile of the characters depicted in the input image. Different from image-to-SVG, the model does not reconstruct, but generates a specific character SVG for the input image (see Fig. 6).
We evaluate the alignment between input character images and generated SVGs by prompting GPT-4o [20] to generate a score ranging from 1 to 10, the higher the better. For evaluation metrics, we also calculate the average token length (# tokens) of a generated SVG sample utilizing the Qwen2.5-VL [1] tokenizer. For computational cost, we compute the average generation time of generating per SVG sample.
@@ 4 OmniSVG
SVG Tokenizer. As illustrated in Sec. 3, our MMSVG-2M dataset simplifies an SVG by removing all attributes and using five basic path commands (see Tab. 2). After the simplification, an SVG script $G$ is represented as the combination of $M$ paths, $G=\{P_{i}\}_{i=1}^{M}$. Here, $P_{i}$ is the $i$-th path containing $N_{i}$ commands, $P_{i}=\{C_{i}^{j}\}_{j=1}^{N_{i}}$, where $C_{i}^{j}$ is the $j$-th command in the $i$-th path. Each command is represented as $C_{i}^{j}=(U_{i}^{j},V_{i}^{j})$, containing both the type identifier $U_{i}^{j}\in\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z}\}$ and the corresponding location argument $V_{i}^{j}$. To generate colored SVG contents, we assign special tokens for hex values to control the “Fill” ($\mathrm{F}$) attribute, distinguishing it from the original SVG commands and coordinates. To this end, we are able to use a total six types of commands $U_{i}^{j}\in\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z},\mathrm{F}\}$ to parameterize a colored SVG parameterization. Specifically, our SVG tokenizer transforms SVG scripts $X_{s}$ into an ordered SVG token sequence within the same representation space as the pre-trained VLM. Following IconShop [55], we flatten the layered structure of the SVG script by concatenating different paths into a single command sequence, where each path begins with the drawing commands followed by point coordinates. Therefore, each SVG sequence could be represented as a flattened sequence. As the generation identifier, we apply special tokens like <SOP> and <EOS> to the two ends of a SVG sequence, identifying the begining and ending of a SVG sequence. We assign special tokens for each command type, i.e. $\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z},\mathrm{F}\}$. To shorten the length of the SVG sequence, we further merge the point coordinates into one token with a mapping function: $<x,y>\rightarrow x\times w+y$, where $w$ is the width of the image. The SVG sequence are then lifted into the same embedding space as the pre-trained VLM with a learnable embedding layer.
Model Architecture. OmniSVG leverages the Qwen2.5-VL architecture, a vision-language models (VLMs), which excels in processing both visual and textual inputs, enabling the generation of precise and compact SVG outputs. OmniSVG is trained for text-to-SVG, image-to-SVG, and character-reference SVG generation task. As depicted in Fig. 2, We first tokenize and embed the interleaved text and image inputs as prefix tokens. Then, we tokenize the SVG scripts into a sequence and concatenate it to the end of the prefix token with the SVG tokenizer. The full sequence is then served as the input of the decoder-only language model.
@@ 6 Conclusions
Limitations and Future Work. During inference, OmniSVG generates tens of thousands of tokens for complex samples, which inevitably leads to a considerable generation time. OmniSVG is only bounded by vector style image prompt and fails on natural images as illustrated in Fig. 9. As for future work, recent endeavors on multi-token prediction [15, 2] and KV-cache compression [68, 3] provide a promising way to save the generation cost. Additionally, the auto-regressive nature of OmniSVG also unlocks future opportunities for in-context learning [67, 69, 46], chain-of-thought reasoning [54, 16], and multi-turn interleaved generation [19, 30], thereby providing a more dexterous user control. Moreover, more complicated tasks like stylization [65, 70] is also worth exploring in the future.
@@ A.2 SVG-Image-Text Pairs Construction
Our MMSVG-2M dataset comprises two million SVG samples with the corresponding rasterized images. Following the recent dataset construction pipeline [28], we generate captions on the rasterized images with BLIP-2 [26], thereby providing textual descriptions that enable us to fine-tune our model to follow these instructions. We use CairoSVG [23] for rasterization and remove samples that produced completely white images.
@@ A.3 Character-SVG Pairs Construction
As illustrated in the Fig. 1, part of our proposed MMSVG-2M-Character subset is constructed using a generative pipeline. As shown in the pipeline diagram in Fig. 2, we employ a FLUX [24]-based generative model enhanced with a vector-style LoRA to enable the generation of SVG-style data. For image-based conditioning, we adopt FLUX-Redux [25], which injects image features via a SigLIP encoder and projects them into image embeddings. These embeddings are then concatenated with the text tokens as conditioning inputs for FLUX [24]. However, in practice, the original Redux [25] conditioning proves to be overly strong. To address this, we adopt a community-implemented variant of Redux that downsamples the image embeddings in 2D space. As observed in our experiments shown in Fig. 11, a downsampling factor between 2×2$\times$2 × and 3×3$\times$3 × yields the most reasonable SVG-style character references. Finally, we employ VTracer [12] to perform near-instant vectorization of the generated images. To construct the MMSVG-2M-Character subset, we first filter $103k$ character instances from the Danbooru [13] dataset and apply the aforementioned pipeline with motion and expression keywords like previous works [9, 8, 35, 64]. We compare the raw FLUX [24] outputs and their vectorized counterparts, retaining only those samples with PSNR and SSIM scores above a certain threshold as valid data.